a coin is weighted so that there is a 59.1% chance of it landing on heads when flipped. the coin is flipped 13 times
find the probability that the number of flips resulting in heads is at least 5 and at most 10
Answer #1
  • Please check attachment  

v Bin In=13, p=0.591) Required Probability = P(5.5*310) = P(x=5) + PIX= 6) + BCX=7) + P(x=8) +P(x=9) + PexFloj = 660.59056-0.